# Impurity bands, line-nodes, and anomalous thermal Hall effect in Weyl superconductors

### v1

@@ V Thermal conductivity at low temperature
The factors $X$ and $Y$ stem from the vertex correction of the nonequilibrium Keldysh Green function and describe the nature of impurity bound states in superconductors [28, 29]. Because only unpaired quasiparticles carry entropy, a finite residual (zero-energy) DOS is a necessary, but not sufficient, condition to obtain a finite THC [see Eq. (26)]. As shown in Fig. 4(a), the THC vanishes both in the Born $(\delta\to 0)$ and unitarity $(\delta\to\frac{\pi}{2})$ limit, and exhibits a peak at intermediate the scattering phase-shift. The key observation is that the thermal Hall signal is enhanced by the effective DOS anisotropy at the Fermi surface, i.e., the finite slope of the DOS as a function of energy near $\epsilon=0$. This slope vanishes in both unitarity and Born limit as shown in Fig. 2, but $(\partial N_{s}/\partial\epsilon)_{\epsilon=0}\neq 0$ is realized in the intermediate range of the scattering phase-shift. As discussed in Sec. III, the quasiclassical method does not capture the intrinsic contribution to the THC, and hence, the THC from Eq. (26) vanishes in both limits. In Sec. VI, we clarify the relation between the extrinsic ATHE and the emergence of the impurity band.
Importantly, even at low temperatures, the extrinsic contribution is comparable to the intrinsic contribution when $\delta\gtrsim 0.75\pi$ ($\delta\gtrsim 0.55\pi$) in the chiral $p(f)$-wave state [see Fig. 4(a)].
@@ VI Impurity bound states and anomalous thermal Hall effect
We now demonstrate explicitly how the extrinsic ATHE is closely related to the emergence of the impurity bound states in superconductors. Recall that in unconventional superconductors with a momentum-dependent gap function, multiple Andreev scattering at impurities creates the quasiparticle bound states. This impurity band affects the quasiparticle DOS, especially near the unitarity limit, as seen Fig. 2 [26, 27]. As the phase-shift of impurity scattering deviates from the unitarity, the energy of the individual bound states moves to a finite energy. For weak scatterers, the spectral weight for the bound states is eventually absorbed in the coherence peaks around $\epsilon\simeq\pm|\Delta_{\rm eq}|$.
From the $T$-matrix equation in equilibrium and the zero-energy DOS, we find the relationship between the extrinsic ATHE and the DOS,
@@
Equation (30) is derived in Appendix C. Equation (30) shows that the ATHE is closely related to the scattering phase-shift dependence of the residual DOS, $\frac{\partial n_{s}(0)}{\partial\delta}$, which is characterized by the emergence of the impurity bound states. As seen in Fig. 4(b), the zero-energy DOS rapidly grows in the intermediate scattering phase-shift. The variation of $n_{s}(0)$ with $\delta$ slows down and eventually saturates in the unitarity limit. This evolution of the zero-energy DOS qualitatively agrees with the phase-shift dependence of the ATHE in WSCs.
@@ VII Line-nodes and Anomalous Thermal Hall Effect
We are now in the position to compare the full numerical evaluation results for the thermal transport for the $E_{1u}$ chiral $p$-wave and $f$-wave states, showing the influence of additional line-nodes on the extrinsic ATHE. As predicted by the low temperature expansion analysis, the extrinsic ATHE is small both for weak (Born) and strong (unitarity) scatterers and reaches the largest value for the intermediate phase-shift (see Fig. 5).
The relative magnitude between the intrinsic and extrinsic ATHE is consistent with the result of the low temperature expansion analysis [see Figs. 4(a) and 5]. Assuming that the intrinsic contribution to the THC exhibits $T$-linear and unaltered at higher temperatures, the intrinsic mechanism should dominate over the extrinsic contribution at higher temperatures. However, in that range, we generally expect corrections to the result in Eq. (7).
Notably, the maximum amplitude of the THC in the chiral $f$-wave state with two horizontal line-nodes is smaller than that of the chiral $p$-wave state. This is in sharp contrast with the longitudinal thermal transport which is enhanced by the enlarged phase spaces available to quasiparticles near nodal lines. This result implies that the extrinsic ATHE is weakened by additional line nodes. This observation is consistent with the low temperature analysis discussed in Sec. V. The function $\alpha_{1}(\hat{k}_{z})$ is proportional to $\eta(\hat{k}_{z})$ [see Eqs. (1) and (29)] and changes sign across the additional line-nodes, as illustrated in Fig. 3. It is precisely that this sign change reduces the Fermi surface average of this function and suppresses the extrinsic anomalous thermal Hall response.
@@ VII.2 Coexistence of the chiral $p$-wave and $f$-wave pairings
Figure 6: Anomalous Thermal Hall Effect in Weyl superconductors. (a)Temperature dependence of the extrinsic ATHE in the mixed state of the chiral $p$-wave state and the $E_{1u}$ chiral $f$-wave state. (b) $\Delta_{\rm p}/\Delta_{\rm E1u}$ dependence of the thermal conductivity at low temperature, $T=0.02T_{c,{\rm clean}}$. We set the scattering phase-shift $\delta=\frac{\pi}{12}$(red lines), $\frac{\pi}{6}$(blue lines), $\frac{\pi}{4}$(Green lines), $\frac{\pi}{3}$(orange lines), $\frac{5\pi}{12}$(yellow lines). (c) Schematic picture of the gap structures of the mixed chiral state model. The dotted lines in the panel (a) plot $-\kappa_{xy}^{\rm int}$ using the low temperature formula, where the minus sign of $-\kappa_{xy}^{\rm int}$ is due to compare the relative magnitude with $\kappa_{xy}^{\rm ext}$. The dashed lines in the panel (b) describe $\Delta_{\rm p}/\Delta_{\rm f}$ where two line-nodes merge into one.
Both the chiral $p$-wave and $f$-wave pairings we considered belong to the $E_{1u}$ irreducible representation and hence these are naturally mixed. [76] We denote the chiral $p$ ($f$)-wave order parameter as $\Delta_{\rm p(f)}$ and write the order parameter in the mixed pairing state as,
@@
Figure 7: $\hat{k}_{z}$-dependence of the functions $\alpha_{0}$ and $\alpha_{1}$ for the mixed chiral state model. The red, blue and green curves show the calculated results with $\Delta_{\rm p}/\Delta_{\rm E_{1u}}=0.7,1,1.3$, respectively. We set the scattering rate as $\Gamma_{\rm imp}=0.04\pi T_{c,\rm clean}$ and the phase-shift as $\delta=\frac{\pi}{6}$.
The nodal structure of the resulting gap depends on the ratio $\Delta_{\rm p}/\Delta_{\rm f}$. There are always Weyl nodes at the north and south poles of a spherical Fermi surface. In addition, the superconducting gap has two line-nodes when $-4<\Delta_{\rm p}/\Delta_{\rm f}<1$. These nodes merge at the equator for $\Delta_{\rm p}/\Delta_{\rm E_{1u}}=1$, and, for $\Delta_{\rm p}/\Delta_{\rm f}>1$, the equatorial line-node is lifted, and replaced by a minimum (see Fig. 6 (b)).
As seen in Fig. 6 (b), the extrinsic ATHE is enhanced when the two horizontal line-nodes merge at the equator. When the mixed chiral state possesses a single line-node or small line-node gap minima at the equator, the extrinsic ATHE becomes dominant at low temperatures. When the line-node at the equator splits into two, the extrinsic ATHE is sharply suppressed. As shown in Fig. 6 (c), the small gap minima also suppresses the extrinsic ATHE, but the suppression by the gap minima is more gradual.
This behavior is also understood from the low temperature formula for the extrinsic ATHE (Eq. 26)). As shown in Fig. 7, when we introduce the chiral $p$-wave gap to the chiral $f$-wave state with $10<\Delta_{\rm p}/\Delta_{\rm E_{1u}}<$ 1, two line-nodes get closer to each other. When two line-nodes approach each other, the positive region of $\alpha_{1}(\hat{k}_{z})$ is enlarged, and its momentum average is enhanced. As shown in Fig. 7, when two line-nodes merge at the equator, there is no sign change of the gap function and $\alpha_{1}(\hat{k}_{z})$ becomes a positive function. Absence of the sign change at line-nodes enhances the extrinsic ATHE since now the contribution of additional quasiparticles in the near-nodal regions (appearing in the denominator of $\alpha_{1}$ in Eq. (29)) is not reduced by the sign change in the numerator on performing the Fermi surface average.
@@ VIII Conclusion
In this paper, we addressed the relative magnitude of the intrinsic and extrinsic ATHE in WSCs. The intrinsic ATHE arises from the existence of Weyl nodes and boundary modes and is determined by the structure of the order parameter and the distribution of Weyl nodes in the momentum space. The extrinsic contribution is due to the skew scattering of quasiparticles. We computed the extrinsic contribution to the THC using the quasiclassical Eilenberger formalism. The intrinsic ATHE is suppressed by the small factor, $1/(k_{\rm F}\xi_{0})$, in superconductors, and thus the extrinsic contribution often dominates the thermal Hall responses. The intrinsic contribution relies on the gapless boundary modes and exhibits $T$-linear behavior regardless of the existence of nodal excitation in bulk [79]. Thus, the intrinsic contribution may dominate the thermal Hall response at low temperatures when the impurity potential is weak. Within this formalism, we combined the low temperature analysis and the numerical calculations to specifically focus on the effects of the impurity bands and the additional nodal structure on the extrinsic ATHE.
One of our findings is that the variation of the density of states near the Fermi energy is crucial in determining the transverse transport coefficients. We find that impurities with the intermediate potential strength are more effective in generating the thermal Hall signal than those in the weak (Born) or strong (unitarity) scattering limit. We confirmed this by investigating the phase-shift dependence of the DOS and the ATHE, and associated the result with the growth of the impurity band from the individual impurity resonances.
@@
We showed that the additional line nodal structure significantly reduces the extrinsic contribution to the THC when line-nodes involve the sign change of the order parameter. For such order parameters, the angular momentum of the Cooper pairs is not the same as the winding number around the nodes. Consequently, the skew scattering from different parts of the Fermi surface partially compensates. In contrast, when the line nodes are not accompanied by such a sign change, quasiparticle excitations around the gap minima significantly enhance the thermal Hall response.
@@
Our results can be applied to any chiral superconductor even without Weyl nodes. For an example, we note the chiral $d$-wave pairing, $\Delta(\bm{k})\propto\hat{k}_{z}(\hat{k}_{x}\pm i\hat{k}_{y})$ among candidate order parameters for Sr2RuO4 [85]. The chiral $d$-wave order parameter does not realize Weyl nodes because this compound has quasi-two-dimensional Fermi surfaces and the Fermi surface is absent at $k_{x}=k_{y}=0$. However, the chiral $d$-wave superconducting order involves the broken time-reversal and mirror reflection symmetries, allowing the ATHE [22]. Because the chiral $d$-wave order parameter realizes a line node at $k_{z}=0$ with the sign change, the line nodal excitations suppress the extrinsic ATHE. Hence, the intrinsic contribution will be a substantial part of ATHE in Sr2RuO4, but the detailed balance between it and the extrinsic contribution depends on the details of the Fermi surface and other material-specific parameters.
@@
Our findings provide clear evidence that in large classes of candidate materials, there is likely to be a competition between the intrinsic and extrinsic contributions to the ATHE even at the low temperature, and careful (and potentially material-specific) analysis is needed to understand its behavior in the superconducting state. Our work lays the framework and provides the blueprint for such analysis.

### v2

@@ V Thermal conductivity at low temperature
Because only unpaired quasiparticles carry entropy, a finite residual DOS is a necessary, but not sufficient, condition to obtain a finite impurity contribution to the THC [see Eq. (28)]. As shown in Fig. 4 (a), the impurity contribution to the THC vanishes both in the Born $(\delta\to 0)$ and unitarity $(|\delta|\to\pi/2)$ limits and exhibits a peak at the intermediate the scattering phase-shift. In the case of $\Gamma_{\rm imp}=0.04\pi T_{c,{\rm clean}}$, the corresponding values of the phase-shift are $\delta\simeq 0.43\pi$ for the chiral $p$-wave state and $\delta\simeq 0.39\pi$ for the chiral $f$-wave state. The key observation is that as other transverse transport coefficients, the extrinsic ATHE is enhanced by the effective DOS anisotropy at the Fermi surface, i.e., the finite slope of the DOS as a function of energy near $\epsilon=0$ [28, 29]. This slope vanishes in both Born and unitarity limits as shown in Fig. 2, but $(\partial N_{s}/\partial\epsilon)_{\epsilon=0}\neq 0$ is realized in the scattering phase-shifts between these limits. In Sec. VI, we clarify the relation between the extrinsic ATHE and the emergence of the impurity bands.
Figure 4 (a) shows that even at low temperatures, the impurity contribution to the THC becomes larger than the intrinsic contribution when the scattering phase-shift is within a certain range. For $\Gamma_{\rm imp}=0.04\pi T_{c,\rm clean}$, these ranges are $0.382\pi\lesssim\delta\lesssim 0.497\pi$ for the chiral $p$-wave state, and $0.284\pi\lesssim\delta\lesssim 0.490\pi$ for the chiral $f$-wave state. This result suggests an underlying competition between the intrinsic and extrinsic ATHE, even at low temperatures.
@@ VI Impurity bound states and anomalous thermal Hall effect
We now explicitly associate the extrinsic ATHE to the formation of impurity bands in unconventional superconductors. From the $T$-matrix equation in equilibrium and the residual DOS, we find the relationship between the extrinsic ATHE and the residual DOS,
@@
Equation (32) is derived in Appendix C. As the scattering phase-shift increases, the zero-energy DOS begins to grow when the tails in the DOS from the electron-like and hole-like impurity bands touch the Fermi level [see the orange curves Fig. 2 (a-b)]. This increase in $n_{s}(0)$ with $\delta$ eventually saturates in the unitarity limit [see Fig. 4 (b)]. The factor, $\partial n_{s}(0)/\partial\delta$, captures this variation in the residual DOS due to the change in the impurity levels and links the ATHE with the formation of the impurity bands.
@@ VII Line-nodes and Anomalous Thermal Hall Effect
We are now in the position to compare the full numerical results for the thermal conductivity in the $E_{1u}$ chiral $p$-wave and $f$-wave states, evaluating the influence of additional line nodes on the extrinsic ATHE.
Figure 5 (a) shows the temperature dependence of the impurity contribution to the THC. As predicted by the low-temperature expansion analysis, the maximum amplitude of $\kappa_{xy}^{\rm ext}$ is small when the impurity potential strength is close to the Born or unitarity limit [see the red and yellow curves in Fig. 5 (a)], and becomes large at intermediate phase-shifts between these limits. The relative magnitude of the intrinsic and extrinsic ATHE is consistent with the result of the low-temperature expansion analysis, see Figs. 4 (a) and 5 [see the green, blue, and orange curves in Fig. 5 (a)]. If we assume that the intrinsic (topological) contribution to the THC discussed in Sec. III maintains its $T$-linear growth, the intrinsic mechanism may dominate the extrinsic (impurity-induced) contribution at higher temperatures. However, as the temperature increases, corrections to the result in Eq. (7) are generally expected.
Figure 6: Temperature dependence of the impurity contribution to the THC in the chiral $f$-wave states described by Eq. (33). The dotted lines indicate $-\kappa_{xy}^{\rm int}$ using the low-temperature formula with $(k_{\rm F}\xi_{0})^{-1}=0.01$, while the minus sign enables direct comparison with the extrinsic term, $\kappa_{xy}^{\rm ext}$. We set the scattering rate $\Gamma_{\rm imp}=0.04\pi T_{c,{\rm clean}}$ and the scattering phase-shift $\delta=\frac{\pi}{12}$ (red lines), $\frac{\pi}{6}$ (blue lines), $\frac{\pi}{4}$ (green lines), $\frac{\pi}{3}$ (orange lines), $\frac{5\pi}{12}$ (yellow lines).
Notably, the maximum amplitude of $\kappa_{xy}^{\rm ext}$ in the $E_{1u}$ chiral $f$-wave state with two horizontal line-nodes is smaller than that of the $E_{1u}$ chiral $p$-wave state. This is in sharp contrast with the longitudinal thermal transport, which is enhanced by the enlarged phase space available near nodal lines [see Fig. 5 (b)]. This result implies that the extrinsic ATHE is weakened by additional line nodes in the chiral $f$-wave state. This observation is consistent with the low-temperature expansion analysis discussed in Sec. V. The function $\alpha_{1}(\hat{k}_{z})$ is proportional to $\eta(\hat{k}_{z})$ and changes sign across the additional line-nodes, as illustrated in Fig. 3 (b). This sign change reduces the Fermi surface average of $\alpha_{1}(\hat{k}_{z})$ and suppresses the extrinsic anomalous thermal Hall response.
Let us demonstrate that the sign changes at the line nodes suppress the extrinsic ATHE in the $E_{1u}$ chiral $f$-wave state. To this end, we consider the chiral $f$-wave order parameter on the Fermi sphere,
${\bm{d}}(\bm{k}_{\rm F})=\Delta(\hat{k}_{x}+i\hat{k}_{y})|5\hat{k}_{z}^{2}-1|\hat{\bm{z}}.$ (33)
This chiral state has the same quasiparticle population as the $E_{1u}$ chiral $f$-wave state because this state has two-horizontal line-nodes at $\hat{k}_{z}=\pm 1/\sqrt{5}$ and Weyl points at the north and south poles on the Fermi sphere. However, the sign changes at the line nodes are eliminated by taking the absolute value in Eq. (33). Figure 6 shows that the chiral state described by Eq. (33) exhibits a larger ATHE signal than the $E_{1u}$ chiral $p$-wave and $f$-wave states [see Fig. 5 (a)]. This is because $\langle\alpha_{1}(\hat{k}_{z})\rangle_{\rm FS}$ becomes large by taking the absolute value in Eq. (33).
We also note that the $E_{1g}$ chiral $d$-wave state with $\Delta(\bm{k})\propto k_{z}(k_{x}\pm ik_{y})$ has a line node at the equator, so that the extrinsic ATHE is suppressed. In this case, the extrinsic contributions from the northern and southern hemispheres precisely cancel each other, and the intrinsic contribution dominates the thermal Hall response.
@@ VII.2 Coexistence of the chiral $p$-wave and $f$-wave pairings
Figure 7: (a) Schematic picture of the gap structures, (b) temperature dependence of the impurity contribution to the THC, and (c) $\Delta_{\rm p}/\Delta_{\rm f}$ dependence of the thermal conductivity in the mixed pairing states. The dotted lines in the panel (b) plot $-\kappa_{xy}^{\rm int}/(N(\epsilon_{\rm F})v_{\rm F}^{2})$ using the low-temperature formula with $(k_{\rm F}\xi_{0})^{-1}=0.01$, where the minus sign is due to comparing the relative magnitude with $\kappa_{xy}^{\rm ext}$. The dashed lines in the panel (c) represent $\Delta_{\rm p}/\Delta_{\rm f}$ where two line nodes merge into one. In panels (b-c), we set the scattering rate $\Gamma_{\rm imp}=0.04\pi T_{c,{\rm clean}}$ and the scattering phase-shift $\delta=\frac{\pi}{12}$ (red lines), $\frac{\pi}{6}$ (blue lines), $\frac{\pi}{4}$ (green lines), $\frac{\pi}{3}$ (orange lines), $\frac{5\pi}{12}$ (yellow lines). In panel (c), we set the temperature $T=0.02T_{c,{\rm clean}}$.
Both the $E_{1u}$ chiral $p$-wave and $f$-wave pairings we consider belong to the $E_{1u}$ irreducible representation, and hence these are naturally mixed [76]. We denote the chiral $p$ ($f$)-wave order parameter as $\Delta_{\rm p(f)}$ and write the order parameter in the mixed pairing state as,
@@
Figure 8: The Fermi surface function $\alpha_{1}(\hat{k}_{z})$ in the mixed chiral pairing states [see Eq. (31)]. The red and green curves represent the calculated results with $\Delta_{\rm p}/\Delta_{\rm f}=0.7,1,1.3$, respectively. We set the scattering rate $\Gamma_{\rm imp}=0.04\pi T_{c,\rm clean}$ and the phase-shift $\delta=\frac{\pi}{6}$.
The nodal structure of the resulting gap depends on the ratio $\Delta_{\rm p}/\Delta_{\rm f}$. There are always Weyl nodes at the north and south poles of the Fermi sphere. In addition, the superconducting gap has two line-nodes when $-4<\Delta_{\rm p}/\Delta_{\rm f}<1$. These nodes merge at the equator for $\Delta_{\rm p}/\Delta_{\rm f}=1$. For $\Delta_{\rm p}/\Delta_{\rm f}>1$, the equatorial line node is lifted and replaced by a minimum [see Fig. 7 (a)].
Figure 7 (b) shows that the extrinsic ATHE is enhanced when the two horizontal line nodes merge at the equator. When the mixed chiral state possesses a single line-node or small line-node gap minima at the equator, the extrinsic ATHE becomes dominant even at low temperatures. Figure 7 (c) shows that the extrinsic ATHE is sharply suppressed when the line node at the equator splits into two. As is clear from Fig. 7 (c), the behavior of the ATHE for the gap minima is very similar to the one shown in Fig. 6, emphasizing once again that line nodes suppress the extrinsic ATHE, even though they enhance the longitudinal thermal conductivity.
This behavior is also understood from the low-temperature formula for the extrinsic ATHE [Eq. 28)]. As shown in Fig. 8, when we introduce the chiral $p$-wave gap to the chiral $f$-wave state with $10<\Delta_{\rm p}/\Delta_{\rm f}<$ 1, two line nodes get closer to each other. When the two line nodes approach each other, the positive region of $\alpha_{1}(\hat{k}_{z})$ is enlarged, and its momentum average is enhanced. As shown in Fig. 8, when two line nodes merge at the equator, $\alpha_{1}(\hat{k}_{z})$ has no sign change and becomes a positive function. The absence of the sign change at the line nodes enhances the extrinsic ATHE because the contribution of additional quasiparticles in the near-nodal regions (appearing in the denominator of $\alpha_{1}(\hat{k}_{z})$ in Eq. (31)) is not reduced by the sign change in the numerator on performing the Fermi surface average.
@@ VIII Conclusion
In this paper, we addressed the relative magnitude of the intrinsic and extrinsic ATHE in WSCs. The intrinsic ATHE arises from the Berry curvature generated by Weyl points. The intrinsic ATHE is determined by the structure of the order parameter and the configuration of Weyl points in momentum space. On the other hand, the extrinsic contribution arises from the skew scattering of quasiparticles at impurity sites. Hence, the magnitude of the extrinsic ATHE depends on the detail of parameters for impurities, such as the form of the impurity potential, the scattering rate, and the scattering phase-shift.
The intrinsic ATHE is suppressed by the small factor, $(k_{\rm F}\xi_{0})^{-1}$, in superconductors, and thus the extrinsic contribution often dominates the thermal Hall responses. At the same time, the intrinsic contribution relies on the gapless surface Majorana modes, which ensure $T$-linear behavior in low temperatures regardless of the existence of nodal excitation in bulk [79]. Hence, the intrinsic contribution may dominate the thermal Hall response at low temperatures in clean systems.
We computed the impurity contribution to the THC using the quasiclassical transport theory. We combined the low-temperature expansion analysis and the numerical calculations to evaluate the effects of the impurity bands and the additional nodal structure on the extrinsic ATHE. One of our findings is that the variation of the DOS due to impurity bands near the Fermi energy is crucial for the THC. The impurity contribution to the THC vanishes in both the Born and unitarity limits. This contribution becomes significant for the intermediate impurity potential strength between these limits. The impurity contribution to the THC becomes significant when the tails of the impurity bands touch the Fermi level, amplifying the residual DOS. We investigated the phase-shift dependence of the DOS and the ATHE and confirmed this behavior. Furthermore, we derived the formula associating the extrinsic ATHE with the residual DOS [Eq. (32)].
@@
We showed that the additional line-nodal structure significantly reduces the impurity contribution to the THC when line nodes involve the sign change of the order parameter. For such order parameters, the angular momentum of the Cooper pairs is not the same as the winding number around the nodes. Consequently, the skew scattering from different parts of the Fermi surface partially compensates. In contrast, when the line-nodes are not accompanied by such a sign change, quasiparticle excitations around the gap minima significantly enhance the thermal Hall response.
@@
Our results can be applied to any chiral superconductor, even without Weyl nodes. For an example, we note the chiral $d$-wave pairing, $\Delta(\bm{k})\propto\hat{k}_{z}(\hat{k}_{x}\pm i\hat{k}_{y})$ among candidate order parameters for Sr2RuO4 [85]. The chiral $d$-wave order parameter does not generate Weyl points because this compound has quasi-two-dimensional Fermi surfaces, and the Fermi surface is absent at $k_{x}=k_{y}=0$. Although this compound is not a WSC, the chiral $d$-wave superconducting order involves the broken time-reversal and mirror reflection symmetries and allows the ATHE [22]. As described in Sec. VII, this chiral $d$-wave order generates a line-node at $k_{z}=0$ with the sign change of the order parameter, and the extrinsic ATHE is suppressed by the line-nodal excitations. Hence, the intrinsic contribution will be a substantial part of ATHE in Sr2RuO4, but the detailed balance between the intrinsic and extrinsic contributions depends on the details of the Fermi surface and other material-specific parameters.
For multi-band materials, we need to sum the contributions from each band to obtain the total THC. Our result can be used to estimate the contributions from each band. In principle, the interband effects modify the ATHE, but unless the system is fine-tuned so that the Fermi level lies near degeneracy points, interband contributions are suppressed due to the energy mismatch between the bands at a given wave vector. While the detailed analysis of this situation remains a subject for future studies, we expect that our results remain at least qualitatively, and likely quantitatively, valid in most such cases.
@@
Our findings provide clear evidence that in large classes of candidate materials for WSCs, there is likely to be a competition between the intrinsic and extrinsic contributions to the ATHE even at low temperatures. This suggests that careful (and potentially material-specific) analysis is needed to understand its behavior in the superconducting state. Our work lays the framework and provides the blueprint for such analysis.